# LLM2TEA: Agentic AI Designer Finds Innovative Objects with Generative Evolutionary Multitasking

### v2

@@ V-C5 Quantization of Innovative Solutions
We demonstrate the designs discovered with LLM2TEA can be 3D printed. As the text-to-3D generative model (Shape-E) tends to generate designs that are non-conforming to manufacture in real-world conditions, most designs uncovered in the text-to-3D generative model possessed some degree of defects and fragments. For instance, as shown in the top row of Fig. 10, some defects such as rough surfaces, sharp edges, open holes, and weak or thin structures, especially in the internal cavities, can easily be resolved with minor repairing and infills to the designs before 3D printing. On the other hand, the non-flat and intricate surface details require printing orientation in a vertical position with minimal breakaway or tree-like printing supports. In addition, the 3D designs require appropriate scaling to preserve the detailed geometrical features during 3D printing.
Based on these considerations, we employ Off-the-shelf stereolithography (SLA) 3D printing given its performance in retaining such intricate details as well as its suitability for small-quantity prototyping. The resulting 3D printed designs are shown in the bottom row of Fig. 10, closely matched with the digital visualization of the designs. The conceptual design process could be accomplished within weeks, from discovering to 3D printing the designs, with the starting to examine the designs during a single evolutionary run in a matter of hours. This demonstration showcases LLM2TEA capability to transform AI-generated outputs into tangible physical objects, underscoring its potential applications in fast design prototyping.
Concluding Remark: The insights presented in this experiment discussion section will help researchers to understand the challenges in the problem of interest and overcome the mentioned technical generation issues in LLMs and GenAI models with our proposed strategies. There several ways practitioners can setup LLM2TEA framework for their own needs, including defining the task domains with the variable , choosing a preferred physics simulator, and setting the hard constraint bound values and to define the exploration behavior. We learned that it is crucial to identify task domains that are sufficiently general, exhibit positive transfer of geometric characteristics between tasks, and integrate well with automatic mesh generation and simulation. Moreover, generated cross-domain objects should maintain validity for evaluations as exploration goes deeper into out-of-domain regions. In summary, our proposed method is applicable to any applications (including autonomous driving and robotics) as long as the generated designs can be evaluated. These findings have broader applicability to research communities, such as those in product design or animation, that aim to leverage LLM2TEA, LLMs, and GenAI models for discovering new designs.
@@ VI Conclusion and Outlook
In this paper, we introduce LLM2TEA, the first agentic AI designer within a generative evolutionary multitasking (GEM) framework, leveraging large language models (LLMs) to explore and discover innovative, physically viable designs across multiple domains. By integrating language-guided evolutionary search, text-to-3D generation, semantic evaluation, and physics simulation, LLM2TEA facilities designers in navigating vast, cross-domain design spaces, unlocking innovative solutions that ground aesthetic novelty with real-world practicality. LLM2TEA serves as a collaborative AI design partner in the creative process, expanding the boundaries of conventional design. Experimental results demonstrate significant improvements in design diversity (up to 174%) and physical performance (with over 73% surpassing leading baselines). The successful 3D printing of several generated designs further validates LLM2TEA’s real-world impact, underscoring its transformative potential as a catalyst for AI-driven, cross-domain innovation and practical design optimization.
While the empirical studies of the prototypical design optimization problem are an abstraction of real-world engineering design scenarios, the potential of the LLM2TEA in accelerating the conceptual design process and enhancing the accessibility of such solutions across multiple disciplines has been demonstrated. Nevertheless, this research direction is still in its infancy stage, and further explorations are encouraged, such as employing prompt evolution [1], evaluating LLM2TEA effectiveness with real-world scenario simulations, and investigations in other scientific and engineering scenarios, to name a few.

### v3

@@ V-C5 Quantification of Novel Solutions
The designs discovered with LLM2TEA can be 3D printed. Since the text-to-3D generative model (Shape-E) tends to generate designs that are non-conforming to manufacture in real-world conditions, most designs uncovered in the text-to-3D generative model possessed some degree of defects and fragmentation. For instance, as shown in the top row of Fig. 10, common defects, such as rough surfaces, sharp edges, open holes, and weak or thin structures, especially within internal cavities, can be easily resolved with minor repairs and infills to the designs before 3D printing. On the other hand, the non-flat and intricate surface details require printing orientation in a vertical position with minimal breakaway or tree-like printing supports. Additionally, the 3D designs require appropriate scaling to preserve its detailed geometric features during 3D printing.
Taking these factors into account, off-the-shelf SLA 3D printing is used due to its ability to retain such intricate details and its suitability for small-quantity prototyping. The resulting 3D-printed designs are shown in the bottom row of Fig. 10, closely matching the digital visualization of the designs. The conceptual design process can be accomplished within weeks, from discovery to 3D printing the designs, with the start of examining the designs during a single evolutionary run in a matter of hours. This demonstration highlights the capability of LLM2TEA to transform AI-generated design outputs into tangible physical designs, underscoring its potential applications in fast design prototyping.
Concluding Remark: We hope that the insights presented in this experiment discussion section will provide researchers a better understanding of the challenges in the problem of interest and address the identified technical generation issues in LLMs and GenAI models using our proposed strategies. There are several ways practitioners can set up LLM2TEA to suit their specific needs, including defining the task domains with the variable , choosing a preferred physics simulator, and setting the hard constraint bound values and to control the exploration behavior. We learned that it is crucial to identify task domains that are sufficiently general, exhibit positive transfer of geometric characteristics between tasks, and integrate well with automatic mesh generation and simulation. Moreover, generated cross-domain designs should maintain validity for evaluations as exploration goes deeper into out-of-domain regions. In summary, the proposed method applies to any applications in science and engineering as long as the generated designs can be evaluated. These insights hold broader relevance to research communities, such as those in product design or animation, and other fields seeking to leverage LLM2TEA, LLMs, and GenAI models for novel design discovery.
@@ VI Conclusion and Outlook
This paper presents LLM2TEA, the first agentic AI designer operating with GEM, leveraging LLMs to explore and discover novel, physically viable designs across multiple domains. By integrating LLM-guided evolutionary search, text-to-3D generation, semantic evaluation, and physics simulation, LLM2TEA empowers designers to navigate vast, cross-domain design spaces, unlocking novel solutions that ground aesthetic novelty with real-world practicality. In particular, LLM2TEA is an agentic AI designer capable of expanding the boundaries of conventional design by leveraging knowledge across multiple domains. It explores a broader design space that extends beyond individual areas and embraces their overlap through the paradigm of evolutionary multitasking. This approach promotes the discovery of novel cross-domain designs through effective implicit knowledge transfer. Experimental results demonstrate significant improvements in design diversity (up to ) and physical performance (with over surpassing leading baselines). The successful 3D printing of several generated designs further validates the real-world feasibility of LLM2TEA, underscoring its transformative potential for AI-driven, cross-domain innovation and practical design optimization.
While the empirical studies of the prototypical design optimization problem are an abstraction of real-world engineering design scenarios, the potential of the LLM2TEA in accelerating the conceptual design process and enhancing the accessibility of such solutions across multiple disciplines has been demonstrated. Nevertheless, this research direction remains in its infancy stage, and further explorations are encouraged in other scientific and engineering scenarios.
Acknowledgment
Melvin Wong gratefully acknowledges the financial support from Honda Research Institute Europe (HRI-EU). The work is also partially supported by the National Research Foundation (NRF), Singapore, through the AI Singapore Programme under the project titled “AI-based Urban Cooling Technology Development” (Award No. AISG3-TC-2024-014-SGKR), as well as by the College of Computing and Data Science (CCDS) at Nanyang Technological University (NTU). The authors would like to thank Mr. Jason Fow Kow, Mr. Lim Rui Yi Ray, Ms. Sun Sitong, and Ms. Swaminathan Navitraa for their assistance in 3D printing the AI-generated designs and supporting video production. Grammarly and ChatGPT were used as writing assistance tools to enhance the readability of the manuscript.